### v1

# An End-to-End Smart Predict-then-Optimize Framework for Vehicle Relocation Problems in Large-Scale Vehicle Crowd Sensing

Abstract
Ubiquitous mobile devices have catalyzed the development of vehicle crowd sensing (VCS). In particular, vehicle sensing systems show great potential in the flexible acquisition of extensive spatio-temporal urban data through built-in smart sensors under diverse sensing scenarios. However, vehicle systems like taxis often exhibit biased coverage due to the heterogeneous nature of trip requests and varying routes. To achieve a high sensing coverage, a critical challenge lies in how to optimally relocate vehicles to minimize the divergence between the spatio-temporal distributions of vehicles and target sensing distributions. Conventional approaches typically employ a two-stage predict-then-optimize (PTO) process: first predicting real-time vehicle distributions and subsequently generating an optimal relocation strategy based on the predictions. However, this approach can lead to suboptimal decision-making due to the propagation of errors from upstream prediction. To this end, we develop an end-to-end Smart Predict-then-Optimize (SPO) framework by integrating optimization into prediction within the deep learning architecture, and the entire framework is trained by minimizing the task-specific matching divergence rather than the upstream prediction error. Methodologically, we formulate the vehicle relocation problem by quadratic programming (QP) and incorporate a novel unrolling approach based on the Alternating Direction Method of Multipliers (ADMM) within the SPO framework to compute gradients of the QP layer, facilitating backpropagation and gradient-based optimization for end-to-end learning. The effectiveness of the proposed framework is validated using real-world taxi datasets in Hong Kong. Utilizing the alternating differentiation method, the general SPO framework presents a novel concept of addressing decision-making problems with uncertainty, demonstrating significant potential for advancing applications in intelligent transportation systems.
keywords:
Vehicle Relocation , Vehicle Crowd Sensing (VCS) , Smart Predict-then-Optimize (SPO) , Alternating Direction Method of Multipliers (ADMM) , Unrolling Approach
journal: Transportation Research Part C: Emerging Technologies
1 Introduction
In recent years, the development of ubiquitous mobile devices such as mobile phones has significantly advanced the field of mobile crowd sensing (MCS) (Ganti et al. 2011). Compared with the traditional fixed sensor networks, MCS offers more flexible and extensive coverage of sensing information while reducing the installation and collection cost (Li et al. 2018a). Individuals participating in MCS utilize the built-in smart sensors in mobile devices such as cameras and GPS, to sense the surroundings and gather city-wide information (Li et al. 2022b).
The MCS can be categorized into community sensing (CS) and vehicle crowd sensing (VCS) based on different types of sensor hosts (Ji et al. 2023a). CS utilizes sensors installed in mobile devices, while VCS leverages the sensing capabilities of vehicles to collect city-wide data (Xu et al. 2019). Host vehicles of VCS include taxis (Honicky et al. 2008), trams (Saukh et al. 2014), buses (Kang et al. 2016), and unmanned aerial vehicles (UAVs) (Li et al. 2022a). In general, the advantages of VCS in terms of temporal duration, geographical coverage, reliability, and consistency make it particularly suitable for scenarios requiring continuous, full coverage, and high-quality data (Ji et al. 2023a). Consequently, VCS is instrumental in domains that require high-precision data, such as infrastructure monitoring (Song et al. 2020, Hasenfratz et al. 2015, Liu and Ma 2022), traffic state estimation (Guo et al. 2022, Zhu et al. 2014) and traffic management (Ji et al. 2023b, Dai and Han 2023).
Specifically, the vehicles in the VCS framework are classified into two types: dedicated vehicles (DVs) and non-dedicated vehicles (NDVs). DVs, such as vehicles equipped with air pollution sensors and delivery drones with cameras, are fully guided by the dispatching center to prioritize sensing tasks. In contrast, NDVs, such as e-hailing vehicles and taxis, are under the complete control of their drivers. Upon receiving a sensing task, DVs will proceed immediately to the specific sensing locations. In contrast, NDVs operate under the control of drivers and are confined to their preferred cruising areas and routes. Moreover, NDVs can be converted into DVs through incentives such as monetary compensation.
Figure 1 provides an overview of the sensing procedure in the VCS framework. The data request platform initially publishes sensing tasks. Upon receiving these tasks, the dispatching center develops relocation strategies for DVs. During the subsequent time intervals, all the sensing vehicles, including DVs and NDVs, move to the specified destinations to perform the assigned sensing tasks. After collecting the necessary data, the sensing vehicles will transmit sensing information back to the data request platform.
Figure 1: An illustration of the sensing procedure in the VCS framework.
In the domain of VCS, the target distribution of sensing tasks differs markedly from the spatio-temporal distribution of sensors. Target distributions typically vary based on specific sensing objectives and adapt over time and space. For instance, monitoring air quality may necessitate uniformly distributed data across a city (Bales et al. 2012, Chen et al. 2018), whereas tasks focusing on industrial pollution may require the spatial distribution in Gaussian (Khedo et al. 2010, Paulos et al. 2007). In contrast, the spatio-temporal distributions of sensors are heavily influenced by the movements of host sensing vehicles (Anjomshoaa et al. 2018).
Therefore, a critical challenge in VCS lies in the effective relocation of vehicles to align with the target sensing distribution. To be specific, the dispatching center needs to develop relocation strategies for DVs to minimize the divergence between the target sensing distributions and vehicle distributions. The state-of-the-art approach to addressing the vehicle relocation problem in VCS is a two-stage predict-then-optimize (PTO) method, where a prediction model is first established to forecast the movements of NDVs (Chang et al. 2023, Simchi-Levi 2014, Mišić and Perakis 2020). The prediction is then used as the input for the downstream optimization problem related to DVs. However, a key limitation of the PTO approach is that training the prediction model is mainly based on empirical error minimization, and metrics such as Mean Squared Error (MSE) and Mean Absolute Error (MAE) are used. The trained prediction model may not necessarily yield optimal strategies in the subsequent optimization stage, because the focus of the optimization stage is task-specific, which may not align well with the metrics of MSE or MAE (Bengio 1997, Ford et al. 2015, Elmachtoub and Grigas 2022, Yan and Wang 2022). This discrepancy highlights the need for more integrated approaches that bridge prediction and optimization phases coherently in the context of VCS.
Recently, the end-to-end Smart Predict-then-Optimize (SPO) framework has been proposed and attracted wide attention (Elmachtoub and Grigas 2022). This framework integrates prediction and optimization within a unified deep-learning architecture. Specifically, the optimization module operates as a differentiation layer, mapping inputs directly to optimal solutions without requiring explicit closed-form solutions. This approach enables the entire neural network to be trained by directly minimizing task-specific errors. However, the application of the SPO framework to address challenges such as vehicle relocation has not been extensively explored.
The main challenge in applying the SPO framework is integrating the optimization layer seamlessly within the deep learning architecture. A widely used approach is the implicit differentiation, where the Karush-Kuhn-Tucker (KKT) conditions of the optimization problems are differentiated to obtain first-order derivatives of the optimal solution with respect to the parameters. However, this method can be computationally demanding in directly computing the Jacobian matrix of the KKT conditions for large-scale problems (Diamond and Boyd 2016, Agrawal et al. 2019, Tang and Khalil 2022).
In this paper, we advocate an alternative approach - explicit unrolling, which avoid the direct computation of the Jacobian matrix in the KKT conditions and compute the corresponding gradients iteratively, enhancing computational efficiency and suitability for large-scale problems (Sun et al. 2023, Li et al. 2020). Nevertheless, the majority of unrolling approaches are predominantly applicable to unconstrained problems. In the task of VCS, relocation challenges are usually formulated as constrained optimization problems. To this end, we incorporate the Alternating Direction Method of Multipliers (ADMM) to solve constrained optimization problems by decomposing the original problem into a series of unconstrained sub-problems (Boyd et al. 2011). Combining the ADMM with differentiation methods and unrolling approaches can potentially address the challenge of integrating large-scale constrained convex optimization problems into deep learning architectures.
To conclude, there are two major issues with the predict-and-optimize problem in VCS. First, the conventional PTO framework for vehicle relocation problems may not generate an optimal relocation strategy, resulting in a substantial divergence between the desired target distribution and the actual vehicle distribution. Second, embedding an optimization layer within neural networks is usually computationally intensive and impractical for large-scale networks, and existing unrolling approaches are often inadequate for handling constrained optimization problems effectively.
To address these challenges, we develop an innovative SPO framework coupled with an alternating differentiation method for vehicle relocation problems in VCS, aiming to achieve optimal sensing coverage. Specifically, we formulate the vehicle relocation problem based on constrained quadratic programming (QP) and embed the QP layer into neural networks. To our knowledge, this is the first time that the SPO framework coupled with an alternating differentiation method has been applied in vehicle relocation in mobile sensing. Our contributions are outlined as follows:
  • 1.
    A novel end-to-end Smart Predict-then-Optimize framework for vehicle relocation problems. We first time formulate the vehicle relocation problem using an SPO framework in deep learning. Specifically, the vehicle relocation problem is formulated as a constrained QP, where the parameters in its objectives are predicted using deep learning.
  • 2.
    An unrolling approach using alternating differentiation methods based on the ADMM for constrained optimization problems in large-scale networks. A novel unrolling approach is proposed to obtain the gradients of the constrained QP iteratively for large-scale networks. Specifically, the forward pass is solved using the ADMM, and the backpropagation is achieved through alternating differentiation.
  • 3.
    Real-world experiments. Numerical experiments on a large-scale network with real-world data are conducted. Superior performance and unique characteristics of the proposed SPO framework are demonstrated.
The remainder of the paper is organized as follows: Section 2 reviews the related work on the vehicle relocation problem in mobile sensing. Section 3 introduces the SPO framework in vehicle relocation, including the prediction module, the optimization module, the integrated model, and an unrolling approach for the SPO framework. Section 4 presents the overall solution algorithm of the end-to-end SPO framework. Section 5 evaluates the proposed framework in real-world scenarios and compares the method with the baseline approaches. Section 6 concludes the results and proposes future directions.
2 Literature Review
This section provides a comprehensive review of the literature related to the vehicle relocation problem in mobile sensing, the end-to-end SPO framework and its differentiation approach, and applications of the SPO framework in the field of transportation.
2.1 Vehicle relocation in mobile sensing
Pervasive crowd sensing plays an important role in intelligent transport systems and smart city applications (Ji et al. 2023a, Dai and Han 2023). Given their relatively lower installation costs and extended operational duration, taxis are the most commonly utilized hosts in the VCS. However, due to heterogeneous trip patterns, the spatio-temporal distributions of these vehicles are often biased. To enhance sensing quality, numerous scholars have investigated intervention measures to effectively relocate vehicles through the development of incentivizing schemes (Fan et al. 2019, Xu et al. 2019, Chen et al. 2020) or proactive scheduling (Masutani 2015).
Existing research on solving the conventional vehicle relocation problem can be categorized into three types: the steady-state method, the model-free method, and the real-time optimization method (Lei et al. 2020, Qian et al. 2022). The steady-state method modeled the vehicle relocation system through a queuing theoretical framework and solved the optimal relocation strategy when the system reaches a steady state (Pavone et al. 2012, Sayarshad and Chow 2017, Chuah et al. 2018). A significant limitation of this approach is the inadequate capacity to manage the complexities of large-scale problems. In contrast, the model-free method, based on deep reinforcement learning (DRL), has emerged as a promising approach for its merits in addressing large-scale vehicle relocation challenges (Qian et al. 2022, Jiao et al. 2021, Shou and Di 2020, Qin et al. 2022).
Another classical approach in vehicle sensing is the real-time optimization method, which formulates the vehicle relocation problem from an optimization perspective and identifies optimal strategies through various solutions (Kek et al. 2009, Huang et al. 2018). Miao et al. (2016) and Zhang et al. (2016) formulated the vehicle scheduling and routing problem as a mixed integer program (MIP) and utilized the model predictive control (MPC) method to solve it. Recent advancements have also incorporated deep learning techniques to forecast future travel patterns, employing a two-stage PTO framework for the vehicle relocation problem (Weikl and Bogenberger 2013, Chang et al. 2022; 2023). This method first estimates the spatio-temporal vehicle patterns using deep learning, followed by solving the downstream optimization problem with the prediction results as inputs. However, a notable limitation of the PTO approach is that training the prediction model solely based on prediction error can result in suboptimal decision-making, as opposed to directly minimizing the decision error (Agrawal et al. 2019, Elmachtoub and Grigas 2022). Even with optimal predictions yielding minimal prediction errors, the resulting optimization may not be optimal. Our study extends the real-time PTO framework by integrating optimization directly within the prediction process through a deep learning architecture.
Table 1 summarizes related works of the vehicle relocation problem in mobile sensing and their fundamental settings. In a nutshell, our work differs from the previous research in three key aspects. First, unlike most existing literature that fully controls vehicles, we classify vehicles into dedicated and non-dedicated types, allowing the non-dedicated ones to operate freely. Second, to our knowledge, we are the first to apply the SPO framework to vehicle relocation in mobile sensing and validate it on large-scale networks. Third, we evaluate the framework across multiple target distributions, an aspect rarely explored in previous studies.
Table 1: The related work of vehicle relocation in mobile sensing.
References Fleet type Control ratio Target distribution Solution approach
Miao et al. (2016) Taxi Fully Reality MPC
Xu et al. (2019) Taxi Partially U, G, GM MPC
Wang et al. (2019) Not specific Partially Random PTO
Chen et al. (2020) Taxi Partially Balanced DRL
Fan et al. (2021) Taxi, DV Fully N,U OPT
Jiao et al. (2021) Ride-hailing Fully - DRL
Xu et al. (2021) UAV Fully R OPT
Qian et al. (2022) Taxi Fully U DRL
Chang et al. (2023) Bike Fully Predicted PTO
Jiang et al. (2023) Taxi Fully U OPT
Ours Taxi Partially U, G, GM SPO
  • 1.
    U: Uniform distribution; G: Gaussian distribution; GM: Gaussian mixture distribution; N: Normal distribution; OPT: optimization without prediction.
2.2 The end-to-end SPO framework and its differentiation methods
The SPO framework is an emerging paradigm in the data-driven optimization field that can leverage deep learning tools with potential applications in inventory management and electric grid scheduling etc. (Elmachtoub and Grigas 2022, Donti et al. 2017). Unlike traditional neural networks that utilize explicit formulations, the SPO framework integrates an optimization module as an implicit differentiation layer. This approach maps the inputs of the implicit functions within the optimization layer to optimal solutions. Consequently, training the optimization layer is not easy due to the absence of explicit closed-form derivatives.
To compute the gradients of the optimization layer, two principal methods are employed: the implicit differentiation method and the explicit unrolling approach. The implicit method involves calculating the first-order derivatives of the optimal solution concerning the parameters, typically by directly differentiating the KKT conditions. Tools such as CVXPY (Diamond and Boyd 2016), Cvxpylayer (Agrawal et al. 2019), PyEPO (Tang and Khalil 2022) and OptNet (Amos and Kolter 2017) facilitate this process for various problem types, including linear programming (LP), integer programming (IP) and QP. However, implicit differentiation methods necessitate extensive computation of the Jacobian matrix, rendering them unsuitable for large-scale network problems (Sun et al. 2023). An alternative approach is the explicit unrolling approach, which applies an iterative first-order gradient method. Existing research on explicit unrolling approaches is limited. A notable contribution is by Sun et al. (2023), which proposed an alternating differentiation method based on the ADMM to differentiate convex optimization problems with polyhedral constraints efficiently and recursively. The alternating differentiation method decouples the optimization layers in an alternative manner, thereby reducing the dimension of the Jacobian Matrix. Li et al. (2020) proposed both implicit and explicit methods for differentiating a Variational Inequalities (VIs) layer with a fixed point method. The implicit method leverages the sensitivity of the solutions to the VIs, while the explicit approach unrolls the projection method to solve the VIs.
In summary, implicit differentiation methods directly obtain derivatives but entail significant computational costs for large-scale problems, making them more suitable for simpler optimization problems, such as LP with few constraints. Conversely, explicit differentiation methods unroll the differentiation procedure, reducing the dimension of the inverse Jacobian matrix, and can be more suitable for large-scale and complex convex optimization problems, such as QP.
2.3 Applications of the SPO framework
Recent studies have increasingly investigated the application of the end-to-end SPO framework to traffic operations and management (Yan et al. 2021), in the fields of maritime transportation (Yan et al. 2020, Tian et al. 2023, Yan et al. 2023, Yang et al. 2024) and logistics (Liu et al. 2021, Qi et al. 2023). Yan et al. (2020) developed a semi-SPO framework for efficient ship inspections by utilizing a surrogate mean squared difference in overestimating the number of deficiencies, rather than minimizing the mean squared error. This approach involved first predicting deficiency numbers for each inspector per ship and then constructing an integer optimization model to allocate inspectors to deficient ships. It is noteworthy that in their semi-SPO framework, prediction and optimization remain distinct processes. Similarly, Yang et al. (2024) follows the SPO criterion and proposes a task-specific metric named cumulative detected deficiency number (CDDN) to evaluate the efficiency and effectiveness of port state control inspection. However, the above studies focus on proposing novel SPO metrics rather than embedding an optimization layer in the neural network.
Although the SPO framework has been applied to various fields (Yan et al. 2020, Liu et al. 2021, Yan et al. 2021, Tian et al. 2023, Yan et al. 2023, Qi et al. 2023, Yang et al. 2024, Zhang et al. 2024), its applications in vehicle relocation are overlooked. A closely related study of embedding optimization layer to ours is by Zhang et al. (2024), which applied the end-to-end SPO framework to express pickup and delivery systems. They constructed a Graph Convolutional Network (GCN)-based model to predict order quantities and formulated a K-means clustering model to determine the optimal Areas of Interest (AOI) assignment for couriers. They utilized implicit differentiation by directly differentiating the KKT conditions to enable backpropagation within the neural network.
Table 2: Comparison of the related work in the SPO framework in transportation within the deep learning architecture.
References Optimization layer Constraints Differentiation method Dimension Scalability
Zhang et al. (2024) LP $\surd$ Implicit by KKT 35, 100 Small
Ours QP $\surd$ Explicit by alternating differentiation $2025\ (45\times 45),$ $4624\ (68\times 68)$ Large
Table 2 presents a comparison of studies employing the SPO framework in transportation within the deep learning architecture. Notably, compared to the study by Zhang et al. (2024), our work demonstrates increased complexity in the formulation of the QP layer characterized by a higher dimensionality of decision variables and distinctively employs the alternating differentiation method to address large-scale problems.
3 The SPO framework
This section elaborates on the proposed SPO framework and the embedded modules. We first describe the design of the prediction module, followed by the formulation of the vehicle relocation problem. Subsequently, we introduce the integrated SPO framework and conclude by deriving the solution algorithm for the optimization layer based on the alternating differentiation method.
Figure 2 illustrates the overall framework of the end-to-end SPO vehicle sensing problem. The framework comprises two fundamental modules: a prediction module and an optimization module. The prediction module employs a spatial-temporal neural network designed to forecast the short-term demand of DVs. Additionally, the optimization module formulates the relocation problem of NDVs as a QP model, which generates the optimal relocation strategy and integrates the alternating differentiation method to facilitate backpropagation. The ultimate goal of the SPO framework is to minimize the divergence between the target distribution and the vehicle distribution.
Figure 2: The end-to-end SPO framework of the vehicle sensing problem.
As previously discussed, DVs in the VCS system adhere to designated routes under the direction of the dispatching center, whereas NDVs independently select their cruising areas and pathways. Consequently, when the dispatching center receives sensing requests from the upstream platform, it lacks prior knowledge regarding the future spatial distribution of NDVs, despite their considerable impact on the overall vehicle distribution. To devise an effective relocation strategy for DVs, it is crucial to predict the distribution of NDVs in advance. This predicted demand is then incorporated into the optimization module as a parameter.
3.1 Vehicle demand prediction model for NDVs
This section first introduces the spatio-temporal neural network, then presents the design of the prediction model for NDVs in the SPO framework.
3.1.1 Spatio-temporal vehicle movement prediction
The vehicle demand varies dynamically over time and space. A common approach to manage this variability is discretizing the study area and time length into small intervals for easy implementation of specialized spatio-temporal deep learning models (Ke et al. 2021).
In this paper, the vehicle prediction problem is modeled on a spatio-temporal scale. For the spatial dimension, the study area is partitioned into various regular hexagon grids based on geographical information (Ke et al. 2018). For the temporal dimension, each day is uniformly divided into equal time intervals. Consider a spatial network $\mathcal{G=(V,E)}$, where $\mathcal{V}$ denotes the set of vertices consisting of the set of origin grids $I$ and the set of destination grids $J$, i.e., $\mathcal{V=I\cup J}$, such that $|I|=|J|=N$. The set of edges $\mathcal{E}$ edges connects all adjacent grids.
Let $V$ denote the vehicle class, with each class $v$ defined as $v={c,f,a}$, where $c$ represents controllable DVs, $f$ represents free NDVs, and $a$ encompasses all vehicles. The set of time intervals is denoted as $T$, with each interval of length $\delta$ for each time interval $\tau\in T$. For each time interval $\tau$, the spatial distribution of each vehicle type $v$ is defined as $\mathbf{D}_{v}^{\tau}\in\mathbb{R}^{N}$. To be precise, $\mathbf{D}_{v}^{\tau}=[D_{v,1}^{\tau},...,D_{v,i}^{\tau};...,D_{v,I}^{\tau}]^{\prime}\in\mathbb{R}^{N}$, where $D_{v,i}^{\tau}$ represents the vehicle demand of vehicle type $v$ in grid $i$ at time $\tau$, and is the transpose of a matrix. Then the matching distribution of all vehicles at time $\tau$ is represented by $\mathbf{D}_{a}^{\tau}=\mathbf{D}_{f}^{\tau}+\mathbf{D}_{c}^{\tau}$.
Based on the above notations, the short-term demand prediction model of NDVs is defined as follows: Given the historical observations of NDVs over the past $m$ time intervals $[D_{f}^{\tau-m+1},\ldots,D_{f}^{\tau}]$, the goal is to predict the spatial demand $\mathbf{D}_{f}^{\tau+1}$ for the next time interval $\tau+1$.
3.1.2 TGCN-based prediction model for NDVs
The Temporal Graph Convolutional Network (TGCN) (Zhao et al. 2019), which integrates the Graph Convolutional Network (GCN) and the Gated Recurrent Unit (GRU), is capable of concurrently capturing the inherent spatial dependencies of vehicle demand in adjacent regions and the temporal dynamics. Accordingly, we develop a two-layer prediction model based on the TGCN architecture to forecast the real-time demand of NDVs. The model incorporates a Rectified Linear Unit (ReLU) activation function and is defined as follows:
$\hat{D}_{f,i}^{\tau+1}=\text{ReLU}(\text{TGCN}(\text{ReLU}(\text{TGCN}(D_{f,i}^{\leq\tau},A_{f,i}^{\leq\tau},\hat{H}_{f,i}^{\leq\tau})))),$ (1)
where the $\hat{D}_{f,i}^{\tau+1}$ is the predicted demand of NDVs at time $\tau+1$, which is also the output of the prediction module, $A_{f,i}^{\leq\tau}$ is the adjacency information, $\hat{H}_{f,i}^{\tau}$ is the hidden information. The hat symbol, $\hat{\cdot}$, indicates the variable is an estimator for the true(unknown) variable. $\text{ReLU}(\cdot)=max(0,\cdot)$.
We define the prediction loss in the prediction module in Eq. (2):
$\mathscr{L}_{1}=\lVert{\mathbf{D}_{f}^{\tau+1}-\hat{\mathbf{D}}_{f}^{\tau+1}}\rVert_{2}^{2},$ (2)
where $\mathbf{D}_{f}^{\tau+1}$ is the actual distribution of the NDVs at time $\tau+1$.
Note that any spatio-temporal prediction model can be used in the SPO framework, here we employ the classical TGCN model as an example.
3.2 Vehicle relocation model for DVs
This section initially discusses the variable relationship between prediction and optimization modules, then presents the formulation of the vehicle relocation problem and its vectorized representation.
3.2.1 Connection between prediction and optimization variables
Before proposing the optimization model, we first illustrate the connection among the various variables between prediction and optimization modules, as shown in Figure 3. The output of the prediction layer $\hat{\mathbf{D}}_{f}^{\tau+1}$ is input in the optimization layer as a known parameter.
Figure 3: The variable relationship in different modules.
After predicting the movements of NDVs, we can easily obtain the required (estimated) distribution for DVs by a simple subtraction of the given target distribution $\mathbf{T}_{a}^{\tau+1}$ and the predicted distribution $\hat{\mathbf{D}}_{f}^{\tau+1}$ in Eq. (3):
$\hat{\mathbf{D}}_{c}^{\tau+1}=\mathbf{T}_{a}^{\tau+1}-\hat{\mathbf{D}}_{f}^{\tau+1},$ (3)
where $\mathbf{T}_{a}^{\tau+1}$ is the target distribution for all vehicles at time $\tau+1$.
3.2.2 Formulation of the vehicle relocation model
This section presents the QP formulation of the relocation model for DVs within the optimization module. We regard the original distribution of DVs $\mathbf{D}_{c}^{\tau}$ as supply and the required distribution of DVs $\mathbf{\hat{D}}_{c}^{\tau+1}$ as demand for simplicity. The goal in the subsequent optimization module is to develop an optimal vehicle relocation strategy that satisfies the future demand with the current supply by considering the total incentive cost budget, accessibility constraints, and supply limitations.
Denote $x_{c,ij}^{\tau}\in\mathbb{R}^{N\times N}$ as the dedicated vehicle flow from grid $i$ to grid $j$ during $\tau$ to $\tau+1$. Then the vehicle relocation problem is formulated in Eq. (3.2.2):
{mini!}
[…]
3.2.3 Vectorizing the relocation formulation
To seamlessly integrate the optimization module into the neural network, all two-dimensional variables and parameters will be flattened and vectored into one dimension. Specifically, $x_{c,ij}^{\tau}\in\mathbb{R}^{N\times N}$, $w_{c,ij}^{\tau}\in\mathbb{R}^{N\times N}$ are converted to $\mathbf{y}^{\tau}\in\mathbb{R}^{N^{2}}$, $\mathbf{C}^{\tau}\in\mathbb{R}^{N^{2}}$, respectively. For simplicity, we omit the superscript $\tau$, therefore $\mathbf{y}^{\tau}\in\mathbb{R}^{N^{2}}$, $\mathbf{C}^{\tau}\in\mathbb{R}^{N^{2}}$ is replaced by $\mathbf{y}$, $\mathbf{C}$, as is shown below:
$\mathbf{y}=\left(\begin{array}[]{c}y_{1}\\ y_{2}\\ \vdots\\ y_{N}\\ \hline\cr y_{N+1}\\ y_{N+2}\\ \vdots\\ y_{2N}\\ \hline\cr\vdots\\ \hline\cr y_{N^{2}-N+1}\\ \vdots\\ y_{N^{2}}\end{array}\right)_{N^{2}}=\left(\begin{array}[]{c}x^{\tau}_{11}\\ x^{\tau}_{12}\\ \vdots\\ x^{\tau}_{1N}\\ \hline\cr x^{\tau}_{21}\\ x^{\tau}_{22}\\ \vdots\\ x^{\tau}_{2N}\\ \hline\cr\vdots\\ \hline\cr x^{\tau}_{N1}\\ \vdots\\ x^{\tau}_{NN}\end{array}\right)_{N^{2}},\mathbf{C}=\left(\begin{array}[]{c}w^{\tau}_{11}\\ w^{\tau}_{12}\\ \vdots\\ w^{\tau}_{1N}\\ \hline\cr w^{\tau}_{21}\\ w^{\tau}_{22}\\ \vdots\\ w^{\tau}_{2N}\\ \hline\cr\vdots\\ \hline\cr w^{\tau}_{N1}\\ \vdots\\ w^{\tau}_{NN}\end{array}\right)_{N^{2}}$ (40)
The vehicle relocation problem in Eq. (3.2.2) is then reformulated into a general standardized quadratic form, as presented in Eq. (3.2.3).
{mini}
[…]
$\mathbf{G_{3}}=\mathbf{C}^{\prime}$, $\mathbf{G_{4}}=-\boldsymbol{I_{N^{2}}}$, $\mathbf{h_{1}}=\mathbf{{D}_{c}^{\tau}}$, $\mathbf{h_{2}}=\boldsymbol{0}$, $\mathbf{h_{3}}=\mathbf{R}$, $\mathbf{h_{4}}=\boldsymbol{0}$. $\text{diag}(\cdot)$ is the diagonal matrix, $I_{N^{2}}$ is the Identity matrix in dimension $N^{2}$. $\mathbf{A}$ and $\mathbf{B}$ are sparse matrices, and,
$\mathbf{A}=\left(\begin{array}[]{ccc|ccc|c|ccc}1&0\cdots&0&1&0\cdots&0&&1&0&\cdots 0\\ 0&1\cdots&0&0&1\cdots&0&&0&1&\cdots 0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\cdots&\vdots&\vdots&\vdots\\ 0&0\cdots&1&0&0\cdots&1&&0&0&\cdots 1\\ \end{array}\right)_{N\times N^{2}}$, $\mathbf{B}=\left(\begin{array}[]{ccc|ccc|c|ccc}1&1\cdots&1&0&0\cdots&0&&0&0&\cdots 0\\ 0&0\cdots&0&1&1\cdots&1&&0&0&\cdots 0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\cdots&\vdots&\vdots&\vdots\\ 0&0\cdots&0&0&0\cdots&0&&1&1&\cdots 1\\ \end{array}\right)_{N\times N^{2}}$.
The optimization problem in the vectorized form in Eq. (3.2.3), which includes four unequal constraints and a one-dimensional decision variable (tensor), is equivalent to Eq. (3.2.2). The three key properties of Eq. (3.2.3) are as follows:
  • 1.
    The parameter from the prediction layer $\hat{\mathbf{D}_{f}}^{\tau+1}$, is embedded exclusively in the objective function’s term $q$ and does not appear in any of the constraint terms.
  • 2.
    The vectorized form simplifies the practical implementation of embedding the optimization layer within the neural network, allowing for a seamless connection with the prediction layer and the construction of the entire framework.
  • 3.
    By proposing a generalized form with parameters represented as $G_{n}$ and $h_{n}$, we create a more straightforward representation of the optimization problem, facilitating the derivation of the solution algorithm later on.
3.3 The integrated SPO model
By solving the optimization problem in Eq. (3.2.3), we can generate the optimal strategy for DVs. In particular, the actual allocated distribution of DVs $\mathbf{D}_{c}^{\tau+1}$ is then determined by aggregating the number of DVs that arrive in grids at time $\tau+1$:
$\mathbf{D}_{c}^{\tau+1}=\mathbf{Ay}.$ (41)
In the optimization module, we utilize the predicted distribution of NDVs $\hat{\mathbf{D}}_{f}^{\tau+1}$ at time $\tau+1$ to approximate the actual distribution $\mathbf{D}_{f}^{\tau+1}$, aiding in developing the relocation plan. By the time $\tau+1$, NDVs will have reached their destinations, and all DVs will be allocated according to the relocation plan. Consequently, the actual combined matching distribution for all vehicles can be straightforwardly represented by Eq. (42):
$\mathbf{D}_{c}^{\tau+1}+\mathbf{D}_{f}^{\tau+1}=\mathbf{D}_{a}^{\tau+1}.$ (42)
In the end-to-end SPO framework, the ultimate objective is to minimize the matching divergence between the matching distribution $\mathbf{D}_{a}^{\tau+1}$ and target distribution $\mathbf{T}_{a}^{\tau+1}$ of all vehicles. Therefore, the matching loss is defined as follows:
$\mathscr{L}_{2}=\lVert{{\mathbf{T}_{a}^{\tau+1}}-\mathbf{D}_{a}^{\tau+1}}\rVert_{2}^{2}.$ (43)
Compared to $\mathscr{L}_{1}$ in Eq. (2), which is a mid-term prediction loss of NDVs, $\mathscr{L}_{2}$ in Eq. (43) is the matching loss of all vehicles including DVs and NDVs.
From Eq. (42), it is evident that both the prediction error and relocation error contribute to the overall matching divergence. Therefore, the loss function for the SPO framework is formulated by combining the prediction loss in Eq. (2) and the matching loss in Eq. (43):
$\mathscr{L}_{SPO}=w_{1}\mathscr{L}_{1}+w_{2}\mathscr{L}_{2},$ (44)
where $w_{1}$ is the weight for prediction loss for the NDVs, and $w_{2}$ is the weight for the matching loss for all vehicles.
As illustrated in Figure 2, the vehicle relocation problem defined in Eq. (3.2.3) is embedded in the optimization module and connected with the preceding prediction module in Eq. (1) via gradients. This integration allows the prediction module to inform the optimization process through the backpropagation of gradients, facilitating end-to-end training. Consequently, the SPO framework can be formally presented as follows:
{mini}
[…]
Eq. (3.3) has two crucial properties as summarized below:
  • 1.
    This is a sequential decision-making problem where the solution process progresses from addressing the lower-level constraints to solving the upper-level minimization function. It begins with prediction in Eq. (1), followed by optimization in Eq. (3.2.3), and concludes with the calculation of divergence in Eq. (44).
  • 2.
    The overall nested optimization problem contains a sub-constrained optimization problem (3.2.3) within the constraint section, making it challenging to solve directly and efficiently.
Considering these properties, it is essential to develop novel solution methods to effectively address such nested-constrained optimization problems within the deep learning architecture.
3.4 Solution algorithm
In this section, we will first introduce the computational graph in the SPO framework, then propose the explicit unrolling approach of the SPO framework, including the forward pass by the ADMM, and the backward alternating differentiation method.
3.4.1 The computational graph
The computational graph illustrates how the gradients link and propagate in the neural network in the forward and backward pass. Therefore, we first propose the computational graph in the SPO framework, as is shown in Figure 4. The computation of gradients is essential to enable backpropagation (BP) within the neural network. illustrates the forward and backward pass process in the SPO framework.
To solve the SPO framework as formulated in Eq. (3.3), we derive the gradients of the loss function $\mathscr{L}_{SPO}$ w.r.t. the weight $\mathbf{w}_{p}$ in the prediction module, as is shown in Eq. (45):
$\frac{\partial\mathscr{L}_{SPO}}{\partial\mathbf{w}_{p}}$ $=\frac{\partial\mathscr{L}_{SPO}}{\partial\mathscr{L}_{1}}\frac{\partial\mathscr{L}_{1}}{\partial\hat{\mathbf{D}}_{f}^{\tau+1}}\frac{\partial\hat{\mathbf{D}}_{f}^{\tau+1}}{\partial\mathbf{w}_{p}}+\frac{\partial\mathscr{L}_{SPO}}{\partial\mathscr{L}_{2}}\frac{\partial\mathscr{L}_{2}}{\partial\mathbf{D}_{a}^{\tau+1}}\frac{\partial\mathbf{D}_{a}^{\tau+1}}{\partial\mathbf{y}}\frac{\partial\mathbf{y}}{\partial\mathbf{\hat{D}}_{f}^{\tau+1}}\frac{\partial\hat{\mathbf{D}}_{f}^{\tau+1}}{\partial\mathbf{w}_{p}}.$ (45)
Figure 4: The computational graph to illustrate the backpropagation process.
All the derivatives can be computed analytically and automatically by the Autograd package in PyTorch, except for the $\frac{\partial\mathbf{y}}{\partial\hat{\mathbf{D}_{f}}^{\tau+1}}$ in the optimization layer. Hence, the primary challenge is to compute the gradient of the optimization layer to enable forward and backward propagation. To this end, the gradient of the optimization layers necessitates solving the internal optimization problem first and subsequently differentiating the solution with respect to the parameter provided from the preceding layer.
3.4.2 The unrolling K-layer neural network
Then we proceed to introduce how to derive the gradient of the optimal solution with respect to the predefined parameter from the prediction layer, $\frac{\partial\mathbf{y}}{\partial\hat{\mathbf{D}_{f}}^{\tau+1}}$, using an unrolling approach through the alternating differentiation method. Figure 5 illustrates the unrolling alternating differentiation method and compares it with the implicit differentiation method.
In the implicit differentiation method by KKT condition in Figure 5 (a) and (c), the solution for forward is obtained directly, and the corresponding gradient for the backward pass is computed in a single layer. However, obtaining these gradients directly becomes challenging in large-scale networks due to the high computational cost associated with calculating the Jacobian matrix in the KKT condition.
To address this, we propose an unrolling differentiation approach, as shown in Figure 5 (b) and (d). This approach unfolds the one-layer network into $K$ iterations through an alternating updating procedure. In the forward pass, the problem is solved using the ADMM, where the primal, slack, and dual variables are computed iteratively over $K$ iterations until convergence, indicating that the optimal solution is found. During the backward pass, the gradients of the primal, slack, and dual variables are derived through analytical functions. Each iteration in the forward and backward pass is mapped in a single layer. The final gradient in the optimizer for the alternating differentiation method is obtained by combining the gradients from all $K$ layers. Once convergence is achieved, these $K$ layers are stacked together, forming a $K$-layer neural network. Essentially, the one-layer neural network used in the implicit differentiation method is equivalent to the unrolled $K$-layer neural network.
Figure 5: A comparison of the explicit and implicit differentiation method. (a) and (c) show the forward and backward pass through the implicit differentiation method by KKT, (b) and (d) depict the forward and backward pass through the alternating differentiation method.
Then we will introduce the solution algorithm for the forward pass and backward pass of the unrolled alternating differentiation method, respectively.
Forward pass based on the ADMM. The optimization problem in Eq. (3.2.3) is a constrained quadratic problem with four unequal constraints. It can be solved using the ADMM by transforming the original constrained problem into an unconstrained optimization problem. This section derives the forward pass and the optimal solution via the ADMM.
Note that the parameter from the prediction layer $\hat{\mathbf{D}_{f}}^{\tau+1}$ is embedded in the objective function’s term $q(\hat{\mathbf{D}_{f}}^{\tau+1})$. Here we replace $\hat{\mathbf{D}_{f}}^{\tau+1}$ as $\theta$ as simplicity, and $q(\hat{\mathbf{D}_{f}}^{\tau+1})$ is replaced by $q(\theta)$.
Following the procedures of the ADMM, we first form the augmented Lagrange function of the problem in Eq. (3.2.3) with a quadratic penalty term as follows:
${}\max_{\mathbf{\mu}}\min_{\mathbf{y}\geq 0,\mathbf{s}\geq 0}\mathcal{L}(\mathbf{y},\mathbf{s},\mathbf{\mu};\theta)$ $=\frac{1}{2}\mathbf{y}^{\prime}\mathbf{P}\mathbf{y}+\mathbf{q(\theta)}^{\prime}\mathbf{y}+\sum_{n=1}^{4}\langle\mathbf{\mu}_{n},\;\mathbf{G}_{n}\mathbf{y}+\mathbf{s}_{n}-\mathbf{h}_{n}\rangle$ (46)
$+\frac{\rho}{2}(\sum_{n=1}^{4}{\lVert\mathbf{G}_{n}\mathbf{y}+\mathbf{s}_{n}-\mathbf{h}_{n}\rVert}^{2}),$
where $\mathbf{s}=\left\{\mathbf{s}_{n}|n=1,2,3,4\right\}\geq 0$ is the set of non-negative slack variables corresponding to the four constraints in Eq. (3.2.3), $\mathbf{\mu}=\left\{\mathbf{\mu}_{n}|n=1,2,3,4\right\}$ is the set of dual variables of the four constraints, $\rho>0$ is the penalty term, $\theta$ is the parameter from the prediction layer and $\theta=\hat{\mathbf{D}_{f}}^{\tau+1}$.
Then the primary, slack, and dual variables are updated alternatively based on the ADMM.
$\mathbf{y}_{k+1}=\operatorname*{argmin}_{\mathbf{y}}\mathcal{L}(\mathbf{y},\mathbf{s}_{n,k},\mathbf{\mu}_{n,k};\theta)\quad\forall n\in\left\{1,2,3,4\right\},$ (47a)
$\mathbf{s}_{n,k+1}=\operatorname*{argmin}_{\mathbf{s}_{n}\geq 0}\mathcal{L}(\mathbf{y}_{k+1},\mathbf{s}_{n,k},\mathbf{\mu}_{n,k};\theta)\quad\forall n\in\left\{1,2,3,4\right\},$ (47b)
$\mathbf{\mu}_{n,k+1}=\mathbf{\mu}_{n,k}+\rho(\mathbf{G}_{n}\mathbf{y}_{k+1}+\mathbf{s}_{n,k+1}-\mathbf{h}_{n})\quad\forall n\in\left\{1,2,3,4\right\}.$ (47c)
Then we derive the analytical updating solutions for the primal $\mathbf{y}_{k+1}$, slack $\mathbf{s}_{n,k+1}$, and dual variables $\mathbf{\mu}_{n,k+1}$, respectively.
Taking the first and second-order derivative of Eq. (46) w.r.t. $\mathbf{y}$ and we have $\nabla_{\mathbf{y}}\mathcal{L}$ and $\nabla^{2}_{\mathbf{y}}\mathcal{L}$ as follows:
$\nabla_{\mathbf{y}}\mathcal{L}$ $=\mathbf{P}^{\prime}\mathbf{y}+\mathbf{q(\theta)}+\mathbf{G}_{n}^{\prime}\mathbf{\mu}_{n}+\sum_{n=1}^{4}\rho(\mathbf{G}_{n}^{\prime}(\mathbf{G}_{n}\mathbf{y}+\mathbf{s}_{n}-\mathbf{h}_{n})){},$ (48)
$\nabla^{2}_{\mathbf{y}}\mathcal{L}(\mathbf{y}_{k+1})$ $={\mathbf{P}+\sum_{n=1}^{4}\rho\mathbf{G}_{n}^{\prime}\mathbf{G}_{n}}{}.$ (49)
Let Eq. (48) equals 00 and $\mathbf{y}_{k+1}$ in Eq. (47a) is therefore obtained as follows:
$\mathbf{y}_{k+1}=-{(\mathbf{P}+\sum_{n=1}^{4}\rho\mathbf{G}_{n}^{\prime}\mathbf{G}_{n})}^{-1}(\mathbf{q(\theta)}+\sum_{n=1}^{4}\rho\mathbf{G}_{n}^{\prime}(\mathbf{s}_{n,k}-\mathbf{h}_{n})+\sum_{n=1}^{4}\mathbf{G}_{n}^{\prime}\mathbf{\mu}_{n,k}).$ (50)
Let $U=-{(\mathbf{P}+\sum_{n=1}^{4}\rho\mathbf{G}_{n}^{\prime}\mathbf{G}_{n})}^{-1}$ and $U$ is constant. Then $\mathbf{y}_{k+1}$ can be updated by Eq. (51):
$\mathbf{y}_{k+1}$ $=U(\mathbf{q}(\theta)+\sum_{n=1}^{4}\rho(\mathbf{G}_{n}^{\prime}(\mathbf{s}_{n,k}-\mathbf{h}_{n})+\mathbf{G}_{n}^{\prime}\mathbf{\mu}_{n,k})).$ (51)
Taking the first-order derivative of Eq. (46) w.r.t. $\mathbf{s}_{n,k},\forall n\in\left\{1,2,3,4\right\}$ , the slack variables $\mathbf{s}_{n,k+1}$ in Eq. (47b) are substituted by Eq. (52) with a ReLU function:
$\mathbf{s}_{n,k+1}=\mathbf{ReLU}(-\frac{1}{\rho}\mathbf{\mu}_{n,k}-(\mathbf{G}_{n}\mathbf{y}_{k+1}-\mathbf{h}_{n})).$ (52)
We finally obtain the forward solution of primal, slack, and dual variables in Eq. (53c):
$\mathbf{y}_{k+1}=U(\mathbf{q}(\theta)+\sum_{n=1}^{4}\rho\mathbf{G}_{n}^{\prime}(\mathbf{s}_{n,k}-\mathbf{h}_{n})+\sum_{n=1}^{4}\mathbf{G}_{n}^{\prime}\mathbf{\mu}_{n,k}),$ (53a)
$\mathbf{s}_{n,k+1}=\mathbf{ReLU}(-\frac{1}{\rho}\mathbf{\mu}_{n,k}-(\mathbf{G}_{n}\mathbf{y}_{k+1}-\mathbf{h}_{n}))\quad\forall n\in\left\{1,2,3,4\right\},$ (53b)
$\mathbf{\mu}_{n,k+1}=\mathbf{\mu}_{n,k}+\rho(\mathbf{G}_{n}\mathbf{y}_{k+1}+\mathbf{s}_{n,k+1}-\mathbf{h}_{n})\quad\forall n\in\left\{1,2,3,4\right\},$ (53c)
Backward pass by the alternating differentiation method. Based on the forward solving procedures in Eq. (53c), we can then present the backward differentiation algorithm, which is to compute the derivatives of the primal, slack, and dual variables with respect to predefined parameters alternatively to enable backpropagation.
In the end-to-end SPO framework, we address a real-time vehicle relocation problem in Eq. (3.2.3), where the pre-defined parameter (output) from the prediction layer is denoted as $\hat{\mathbf{D}}_{f}^{\tau+1}$. In this section, we regard $\hat{\mathbf{D}}_{f}^{\tau+1}$ as $\theta$ for simplicity, and $\theta$ is in $\mathbf{q(\theta)}$.
Applying the Implicit Function Theorem (Krantz and Parks 2002) to Eq. (48), then the derivative of the solution $\mathbf{y}^{k+1}$ w.r.t. the parameter $\theta$ can be formulated as:
$\frac{\partial{\mathbf{y}^{k+1}}}{\partial{\theta}}$ $=-\nabla^{2}_{\mathbf{y}}\mathcal{L}(\mathbf{y}_{k+1})^{-1}{\nabla_{\mathbf{y},\theta}\mathcal{L}(\mathbf{y}_{k+1})}=U(\frac{\partial{\mathbf{q}(\theta)}}{\partial\theta}+\sum_{n=1}^{4}\rho\mathbf{G}_{n}^{\prime}\frac{\partial\mathbf{s}_{n,k}}{\partial\theta}+\sum_{n=1}^{4}\mathbf{G}_{n}^{\prime}\frac{\partial\mathbf{\mu}_{n,k}}{\partial\theta}),$ (54)
where the second-order derivative $\nabla^{2}_{\mathbf{y}}\mathcal{L}$ is obtained in Eq. (49).
The derivative of $\mathbf{s}_{n,k+1}$ w.r.t. $\theta$ can be obtained as follows:
$\frac{\partial{\mathbf{s}_{n,k+1}}}{\partial\theta}=-\frac{1}{\rho}\mathbf{sgn}(\mathbf{s}_{n,k+1})\cdot\mathbf{1}^{\prime}\odot(\frac{\partial\mathbf{\mu}_{n,k}}{\partial\theta}+\rho\frac{\partial(G\mathbf{y}_{n,k+1}-\mathbf{h}_{n})}{\partial\theta}).$ (55)
The derivative of $\mathbf{\mu}_{n,k+1}$ w.r.t. $\theta$ can be easily obtained as follows:
$\frac{\partial{\mathbf{\mu}_{n,k+1}}}{\partial\theta}=\frac{\partial\mathbf{\mu}_{n,k}}{\partial\theta}+\rho\frac{\partial(G\mathbf{y}_{n,k+1}+\mathbf{s}_{n,k+1}-\mathbf{h}_{n})}{\partial\theta}.$ (56)
We finally obtain the explicit differentiation function of primal, slack and dual variables as summarized in Eq. (57c):
$\frac{\partial{\mathbf{y}_{k+1}}}{\partial\theta}=U(\frac{\partial{\mathbf{q}(\theta)}}{\partial\theta}+\sum_{n=1}^{4}\rho\mathbf{G}_{n}^{\prime}\frac{\partial\mathbf{s}_{n,k}}{\partial\theta}+\sum_{n=1}^{4}\mathbf{G}_{n}^{\prime}\frac{\partial\mathbf{\mu}_{n,k}}{\partial\theta}),$ (57a)
$\frac{\partial{\mathbf{s}_{n,k+1}}}{\partial\theta}=-\frac{1}{\rho}\mathbf{sgn}(\mathbf{s}_{n,k+1})\cdot\mathbf{1}^{\prime}\odot(\frac{\partial\mathbf{\mu}_{n,k}}{\partial\theta}+\rho\frac{\partial(\mathbf{G}_{n}\mathbf{y}_{k+1}-\mathbf{h}_{n})}{\partial\theta}),$ (57b)
$\frac{\partial{\mathbf{\mu}_{n,k+1}}}{\partial\theta}=\frac{\partial\mathbf{\mu}_{n,k}}{\partial\theta}+\rho\frac{\partial(\mathbf{G}_{n}\mathbf{y}_{k+1}+\mathbf{s}_{n,k+1}-\mathbf{h}_{n})}{\partial\theta},$ (57c)
where $\frac{\partial q}{\partial\theta}=\mathbf{A}^{\prime}$ and $\odot$ is the Hadamard product. In particular, Eq. (57c) is utilized to perform the backward propagation on the computational graph. Note the above forward pass by the ADMM in Eq. (53c) and the backward differentiation in Eq. (57c) is suitable for any quadratic problems with unequal constraints in the form in Eq. (3.2.3).
4 Solution algorithms
The overall solution algorithm of the proposed SPO framework is presented in Algorithm 1, which can be performed in batch form.
Data: Input the historical demand $\mathbf{D}_{f}^{\leq\tau}$, the target distribution $\mathbf{T}_{a}^{\tau+1}$, the control ratio $\gamma$, the total incentive cost budget $R$, the actual travel time matrix $C$, the penalty term $\rho$, the convergence threshold $\xi$.
Result: Output the relocation strategy $\mathbf{y}^{*}$, the matching distribution $\mathbf{D}_{a}$.
1 Initialize Max epoch $N_{e}$, prediction weight $\mathbf{w_{1}}$, matching weight $\mathbf{w_{2}}$ ;
2
3for Epoch $\ e=1,...,N_{e}$ do
4       Input the $\mathbf{D}_{f}^{\leq\tau}$ into prediction module;
5       Predict the distribution of NDVs $\hat{\mathbf{D}}_{f}^{\tau+1}$ in Eq. (1) ;
6       Compute prediction loss $\mathscr{L}_{1}$ in Eq. (2) ;
7       Input $\mathbf{\hat{D}}_{f}^{\tau+1}$ and $\mathbf{D}_{c}^{\tau}$ into optimization module in Eq. (3.2.3);
8       Solve Optimization() ;
9       Aggregate $\mathbf{y}^{\ast}$ and obtain distribution for DVs $\hat{D}_{c}^{\tau+1}$ in Eq. (41) ;
10       Obtain matching distribution $\hat{\mathbf{D}}_{a}^{\tau+1}$ in Eq. (42) ;
11       Compute the SPO loss $\mathscr{L}_{2}$ in Eq. (44);
12       Back propagate and update weights $w_{p}$ in prediction model;
13
14 end for
15
16 Function Optimization():
17       Input the parameter $\theta$ ;
18       Initialize $k\leftarrow 0$ ;
19       Initialize variables and derivatives $\mathbf{y}_{k},s_{i,k},\mathbf{\mu}_{i,k},\frac{\partial{\mathbf{y}_{k}}}{\partial\theta},\frac{\partial{\mathbf{s}_{i,k}}}{\partial\theta},\frac{\partial{\mathbf{\mu_{i}}_{k}}}{\partial\theta}$ ;
20       Initialize parameters $\mathbf{G_{1},G_{2},G_{3},G_{4}},\mathbf{h_{1},h_{2},h_{3},h_{4}},\mathbf{A},\mathbf{B}$, convergence threshold $\xi$ ;
21       while $|\hat{Z}_{k+1}-Z_{k}|\geq\xi$ do
22             Primal update $\mathbf{y}_{k}$ by Eq. (53a), $\frac{\partial{\mathbf{y}_{k}}}{\partial\theta}$ by Eq. (57a) ;
23             Slack update $s_{n,k}$ by Eq. (53b), $\frac{\partial{s_{n,k}}}{\partial\theta}$ by Eq. (57b) ;
24             Dual update $\mathbf{\mu}_{n,k}$ by Eq. (53c), $\frac{\partial{\mathbf{\mu}_{n,k}}}{\partial\theta}$ by Eq. (57c) ;
25             Compute $\hat{Z}_{k}$ ;
26             $k\leftarrow k+1$ ;
27
28       end while
29      return the relocation strategy $\mathbf{y}^{\ast}$ ;
30
31
32
Algorithm 1 Solving procedure for the end-to-end SPO framework.
5 Numerical experiments
In this section, we evaluate the proposed end-to-end SPO framework embedded with an unrolling approach on a real-world taxi dataset in the Kowloon District, Hong Kong SAR. We first introduce the research area, then compare the proposed framework with the SPO framework with an implicit algorithm and conventional two-stage PTO method on mid-size, and large-scale datasets respectively, to verify the effectiveness and efficiency. Then we conduct ablation studies and sensitivity analyses on the impact of various parameters. All the experiments are conducted on a desktop with Intel Core i7-13700K CPU 3.40 GHz $\times$ 32G RAM, 500 GB SSD, GeForce RTX 3090 Ti GPU.
5.1 General settings
5.1.1 The research area
The study area in the Kowloon District, Hong Kong SAR is first discretized into hexagon grids by Uber’s Hexagonal Hierarchical Spatial Index (H3) (Uber H3 2023) at H8 resolution. The average edge length of each hexagon is 531.41 $m$. For the mid-size network, we select 45 grids, and for the large-scale network, 68 grids are chosen, as illustrated in Figure 6.
Figure 6: Overview of the research area in the Kowloon District, Hong Kong SAR.
Real-time taxi information is collected from the HKTaxi application programming interface (API) from March 14 to March 31, 2023, and the taxi demand in each grid is aggregated every 15 minutes. Following preprocessing, the dataset contains a total of 259,530 data samples. The dataset is partitioned into training, validation, and testing sets in a ratio of $8:1:1$.
Figure 7 visualizes the spatio-temporal taxi demand pattern in a week in the research area. From Figure 7 (a), we observe that high-demand areas are concentrated in the west-central and southern regions, particularly around Mong Kok and Hung Hom areas. Grid 62, located in the Mong Kok area, experiences the highest number of taxi requests. Figure 7 (b) and Figure 7 (c) illustrate the daily demand pattern in a week and the hourly demand pattern on Sunday, respectively, for three selected grids and on average. One can see that the temporal patterns vary across different grids with different demands. The average demand ranges from 100 to 200 vehicles per hour. Grid 62 exhibits the highest demand, ranging from 300 to 600 vehicles per hour, with three distinct demand peaks in the morning, noon, and late night. Grid 47 shows diverse peak hours in the morning, noon, and afternoon, occurring at around 8:00, 12:00, and 19:00. Meanwhile, the demand in grid 39 fluctuates smoothly at around 50 vehicles per hour.
Figure 7: The spatio-temporal demand pattern in the research area.
5.1.2 Hyper-parameter setting
The hyper-parameters in the SPO framework are summarized in Table 3 and all parameters in both the prediction and optimization modules are fine-tuned.
Table 3: Configuration of the SPO framework.
Hyper-parameters Values
Optimizer Adagrad
Batch size 64
Learning rate 0.01
Weight decay 0.001
Look-back window 12
Time interval 15 min
Convergence threshold for alternating differentiation method 0.05
penalty term in optimization 2.0
Total travel budget for optimization layer $8,000/10,000/12,000/15,000$
Network scale $45\times 45$ / $68\times 68$
Target distribution Uniform / Gaussian / Gaussian Mixture
5.1.3 Baselines
To compare the importance of both the SPO framework and the embedded alternating differentiation method (SPO-A), we compare the proposed model with three baseline methods: SPO with CVXPY (SPO-C), 2-stage predict-then-Optimize (PTO), and do-nothing method (DON).
  • 1.
    SPO-C: The SPO-C is an end-to-end SPO framework embedded with the implicit differentiation algorithm CVXPY by differentiating the KKT condition.
  • 2.
    PTO: The PTO method is the two-stage predict-then-optimize framework, in which the matching strategy is generated completely based on the prediction result.
  • 3.
    DON: the DON method does not change the original routes of the DVs and lets the drivers drive by themselves.
5.1.4 Target sensing distributions
We create three different target distributions: uniform distribution, Gaussian distribution, and Gaussian Mixture distribution (Xu et al. 2019) based on the one-week data from March 7 to March 13 to represent various potential requirements:
  • 1.
    Uniform distribution. We first calculate the mean distribution of the one-week data in all grids for each time interval, then we sample the uniform data for each time interval in the range $[0.90,1.10]$. Negative values are filtered.
  • 2.
    Gaussian distribution. We sample a Gaussian distribution over the spatial domain based on the mean distribution with a fixed variance of 15 for each time interval. Negative values are filtered.
  • 3.
    Gaussian Mixture distribution. We sample a Gaussian Mixture distribution over the spatial domain based on the mean distribution with two components with fixed variances of 10 and 20 for each time interval. Negative values are filtered.
5.1.5 Evaluation metrics
To evaluate the matching performance of the SPO framework, we adopt two metrics to compare the divergence between matching distribution $\mathbf{D}_{a}^{\tau+1}$ and the target distribution $\mathbf{T}_{a}^{\tau+1}$.
  • 1.
    Root Mean Squared Error (RMSE):
    $\text{RMSE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\mathbf{D}_{a,i}^{\tau+1}-\mathbf{T}_{a,i}^{\tau+1})^{2}}$
  • 2.
    Symmetric Mean Squared Percentage Error (SMAPE):
    $\text{SMAPE}=\frac{100\%}{N}\sum_{i=1}^{N}\frac{|\mathbf{D}_{a,i}^{\tau+1}-\mathbf{T}_{a,i}^{\tau+1}|}{(|\mathbf{D}_{a,i}^{\tau+1}+|\mathbf{T}_{a,i}^{\tau+1}|)/2}$
5.2 Experimental results and comparison
In this section, we compare the performance of the proposed SPO framework with the baseline methods in mid and large-scale networks in dimensions 2,025 $(45\times 45)$ and 4,624 $(68\times 68)$ respectively. The TGCN module is utilized as the prediction module in both the SPO framework and baseline methods. We set different budgets ranging from 8,000 to 15,000 to examine the performance. The ratio of DVs to NDVs is set to $6\colon 4$.
5.2.1 Training curves
Figure 8 illustrates the training curve of the proposed SPO-A framework under Uniform distribution in mid-sized and large-scale networks. Although the training curve exhibits greater fluctuations in large-scale networks compared to mid-sized networks, both configurations converge around Epoch 250.
Figure 8: Training curves of SPO-A under the Uniform distribution in the mid-size and large-scale network (Unit for RMSE: Vehicles/hour).
5.2.2 Matching performance discussion
The performance of the proposed framework and other baseline methods are presented in Table 4.
Table 4: Performance of SPO-A and other baselines in HK dataset under Uniform target distribution (Unit for RMSE: Vehicles/hour).
Grid size Dimension Total budget SPO-A(Ours) SPO-C PTO DON
RMSE SMAPE (%) RMSE SMAPE (%) RMSE SMAPE (%) RMSE SMAPE (%)
$45\times 45$ 2,025 8,000 9.072 31.623 9.510 31.551 9.563 34.081 14.517 52.643
10,000 9.647 31.552 9.541 31.457 9.551 33.987 14.517 52.643
12000 9.589 31.587 9.593 31.558 9.554 31.523 14.517 52.643
15,000 9.059 31.426 9.507 31.431 9.563 34.087 14.517 52.643
$68\times 68$ 4,624 8,000 9.834 33.479 9.962 33.514 10.701 38.166 15.449 56.384
10,000 9.689 32.966 9.820 33.643 10.519 36.955 15.449 56.384
12,000 9.291 30.504 9.678 33.282 10.098 33.162 15.449 56.384
15,000 9.752 33.277 10.096 32.478 10.547 37.495 15.449 56.384
Under the mid-size network with $45\times 45$ grids, the two end-to-end methods, SPO-A and SPO-C show comparable matching performance to the two-stage PTO method. Notably, the SPO-A method outperforms the PTO method under budget constraints of 8,000 and 15,000, showing an improvement of 5.20%. However, in scenarios with a budget of 10,000 and 12,000, SPO-A slightly underperforms compared to SPO-C and PTO by less than 1.15%. Under the large-scale network, one can see that the proposed method consistently outperforms PTO and SPO-C across all budget settings, achieving the highest average matching accuracy. On average, the SPO-A framework improved matching performance by 7.9% compared to PTO and by 2.51% compared to SPO-C across all budget settings.
The results from different network scales indicate that the end-to-end framework presents an obvious enhancement over the two-stage methods under the predict-then-optimize paradigm. Moreover, the proposed SPO with alternating differentiation methods (SPO-A) method shows better and more robust performance in the large-scale network than in the mid-size network, validating the effectiveness of the alternating differentiation method in optimization.
5.3 Ablation study
In this section, we present four groups of ablation models with three different prediction models (TGCN, DCRNN (Li et al. 2018b), and ASTGCN (Guo et al. 2019)) to justify the effect of different prediction and optimization modules: SPO-A, SPO-C, PTO, NOP (no prediction but with the optimization) and DON (no prediction and no optimization). The four grouped models are examined in both mid and large-scale networks under Uniform and Gaussian target sensing distributions. The ratio of DVs to NDVs is set to $6\colon 4$ and the total budget is set to 10,000.
5.3.1 Matching performance discussion
We first discuss the matching accuracy of different prediction models.
The matching results are summarized in Table 5. Overall, the SPO-A methods with different prediction modules achieve robust and satisfactory performance for all scale network sizes and all target distributions in matching accuracy.
Table 5: The comparison of the matching performance of prediction modules in different frameworks under different networks and target distributions (Unit for RMSE: Vehicles/hour).
Prediction module Framework 45-Uniform 45-Gaussian 68-Uniform 68-Gaussian
RMSE SMAPE (%) RMSE SMAPE (%) RMSE SMAPE (%) RMSE SMAPE (%)
TGCN SPO-A (Ours) 9.647 31.552 9.692 36.673 9.689 32.966 10.312 44.287
SPO-C 9.541 31.457 9.858 36.680 9.820 33.643 10.642 46.029
PTO 9.551 33.987 9.704 36.742 10.519 36.955 11.527 48.112
NOP 13.465 47.988 13.581 53.551 14.569 54.450 15.722 52.675
DON 14.517 52.643 19.352 75.164 15.449 56.384 20.583 76.103
DCRNN SPO-A (Ours) 8.779 30.447 8.126 36.368 9.223 35.507 10.186 42.956
SPO-C 8.747 26.317 8.327 36.602 9.710 36.583 10.480 43.236
PTO 8.781 28.895 8.804 40.760 9.710 36.853 11.957 51.171
NOP 13.465 47.988 13.581 53.551 14.569 54.450 15.722 52.675
DON 14.517 52.643 19.352 75.164 15.449 56.384 20.583 76.103
ASTGCN SPO-A (Ours) 8.476 29.591 9.430 46.292 9.225 30.316 10.667 44.606
SPO-C 8.486 29.980 9.496 36.479 9.644 36.360 10.669 44.702
PTO 9.694 30.301 9.986 47.338 9.721 32.610 11.071 45.366
NOP 13.465 47.988 13.581 53.551 14.569 54.450 15.722 52.675
DON 14.517 52.643 19.352 75.164 15.449 56.384 20.583 76.103
  1. (1)
    Justifying the matching performance of different prediction models: When comparing the SPO-A methods with two-stage methods incorporating three similar prediction modules—TGCN, DCRNN, and ASTGCN, we observe that methods utilizing the DCRNN prediction module consistently demonstrate superior matching performance relative to the other two prediction modules. Furthermore, the SPO-A method outperforms other benchmark models in all scenarios except for the Uniform distribution with 45 grids. This indicates the robustness of the SPO framework across various network sizes and target distribution scenarios.
  2. (2)
    Justifying the use of the prediction module: Comparing NOP with SPO-A, we observe a significant decrease in matching performance by 34.01% on average when prediction is removed while retaining optimization. This indicates the crucial role of the prediction module in achieving higher matching performance.
  3. (3)
    Justifying the use of the alternating differentiation method in optimization: Comparing SPO-A and SPO-C, it is evident that substituting the alternating differentiation method with the implicit differentiation method results in a significant decline in matching accuracy in large-scale networks, despite only minor differences between the two frameworks in mid-size networks. This highlights the importance of the alternating differentiation method in large-scale networks.
5.3.2 Prediction performance discussion
We then proceed to compare the prediction accuracy of different prediction modules in SPO-C, SPO-A, and PTO framework and compare the prediction accuracy with matching accuracy under Uniform and Gaussian distribution in mid-size and large-scale networks, as is illustrated in Figure 9.
Figure 9: The prediction accuracy of different prediction models in SPO-A, SPO-C, and PTO framework and the matching accuracy of SPO-A framework under Uniform and Gaussian distribution in mid-size and large-scale networks (Unit for RMSE: Vehicles/hour; Pred., prediction; Match., matching).
  1. (1)
    Justifying the prediction accuracy across different frameworks. For the SPO-A, SPO-C, and PTO frameworks with various prediction models, it is noteworthy that the prediction accuracy of the SPO framework (including both SPO-A and SPO-C) is slightly lower (by 0.2-0.6 in RMSE) than that of the two-step PTO framework in most cases, despite the SPO framework achieving better final matching accuracy. This difference can be attributed to the SPO framework’s training process, which involves a combined prediction and matching divergence, whereas the PTO framework is trained solely based on prediction divergence. This suggests that incorporating matching divergence can reduce the final matching error but may slightly increase the mid-term prediction error.
  2. (2)
    Justifying the prediction accuracy of different prediction models. DCRNN demonstrates the highest prediction accuracy, outperforming TGCN and ASTGCN, with ASTGCN being the least accurate. This trend is also consistent with the matching performance of the three prediction models discussed in Section 5.3.1.
  3. (3)
    Comparing prediction accuracy with matching accuracy. In most cases, the matching RMSE is higher than the prediction RMSE. However, in the Uniform and Gaussian distributions within mid-size networks with 45 grids, we observed that the final matching RMSE is lower than the mid-term prediction RMSE. This is because matching accuracy depends on both prediction accuracy and relocation accuracy, and the relocation RMSE can be relatively lower based on these specific prediction results.
5.4 Sensitivity analysis
This section introduces the two sensitivity analyses of the proposed methods and other baselines on the uncertainty of potential target distributions and control ratios.
5.4.1 Sensitivity analysis on the uncertainty of target sensing distributions
The target distribution may vary randomly for different tasks, so the proposed SPO-A framework should demonstrate robustness across various target sensing distributions. Hence we conduct a sensitivity analysis of the target sensing distribution in Uniform, Gaussian, and Gaussian Mixture distributions every 15 minutes for 24 hours. Note that the Gaussian Mixture distribution diverges most from the original distribution. The ratio of DVs to NDVs is set to $6\colon 4$ and the total budget is set to 10,000. Figure 10 shows the absolute divergence in each grid in all 96 time intervals in a day. Each pixel in each sub-figure represents the absolute value of the divergence between the target distribution and matching distribution. The lighter the color is, the larger the divergence is. Figure 11 depicts the original distribution, matching distribution by SPO-A, and the target distribution at 8:00, 12:00, and 18:00.
Figure 10: Sensitivity analysis of the SPO framework on different time intervals under Uniform, Gaussian, and Gaussian Mixture distributions in large-scale networks.
Figure 11: The spatial comparison of the original, matching, and target distributions in three-time intervals in the large-scale network.
From Figure 10, it is apparent that Grid 62 and Grid 25 have larger divergence especially from 10:00 to 12:00 in all three target distributions. This discrepancy may be attributed to the anticipation of the prediction from free vehicle movements based on weekly patterns. As mentioned in Section 5.1.1, Grid 62 in the Mong Kok area and Grid 25 in the Hunghom area experience a surge in demand during noon, as evident in Figure 11. When the distribution of NDVs deviates from the target distribution, the remaining DVs struggle to compensate for this divergence. Conversely, in grids with smoother demand fluctuations, the matching divergence is less.
Overall, significant fluctuations in actual taxi demand may challenge the ability of the SPO-A framework to effectively satisfy matching distributions. However, if the target distribution diverges substantially from the original distribution, SPO-A is likely to outperform other baseline methods.
5.4.2 Sensitivity analysis on the uncertainty of control ratios
In this section, we analyze how matching performance varies among the SPO-A, SPO-C, and the PTO framework across different control ratios in real traffic networks. Specifically, we examine seven control ratios ranging from 0.2 to 0.8, with increments of 0.1, in both mid and large-scale networks under Uniform, Gaussian, and Gaussian Mixture distributions. We compare the improvement in matching accuracy of the proposed SPO-A framework against the SPO-C and PTO frameworks, using the RMSE of the PTO framework as a reference. The results are displayed in Figure 12.
In the mid-size network, as shown in Figure 12 (a), (b), and (c), with control ratios ranging from 20% to 60%, the SPO-A framework consistently outperforms the SPO-C and PTO frameworks in most cases. The highest improvement of the SPO-A, 7.52%, is observed at a 60% control ratio under the Gaussian Mixture distribution. Additionally, both the SPO-A and SPO-C frameworks demonstrate better matching performance than the PTO framework within this control ratio range. The SPO-C framework performs comparably to the PTO method, showing an improvement of around 2% over the PTO framework.
In the large-scale network, as shown in Figure 12 (d), (e), and (f), the SPO-A framework achieves even better matching performance compared to the mid-size network, with a peak improvement of 13.75% under the Gaussian distribution compared with the PTO framework. The SPO-A method also significantly outperforms both the SPO-C and PTO frameworks, showing more substantial improvements over these two baselines.
However, with control ratios exceeding 60% in mid-size networks and 70% in large-scale networks, the matching performance of the SPO-A framework declines, falling below that of the SPO-C and PTO methods. This decline is attributed to precision errors occurring during the convergence process in the alternating differentiation iterations.
Overall, the SPO-A framework demonstrates significantly better performance than the SPO-C and PTO frameworks within the control ratio range of 0.3 to 0.7. As the control ratio increases, the precision errors in the alternating differentiation method can enlarge, negatively impacting on the final matching performance.
Figure 12: The improvement in the matching performance of SPO-A with SPO-C and PTO frameworks under Uniform, Gaussian, and Gaussian Mixture distribution in mid-size and large-scale networks, using the matching RMSE of the PTO framework as a reference.
6 Conclusions
This paper presents an end-to-end SPO framework coupled with an alternating differentiation method for vehicle relocation problems in mobile sensing, aimed at enhancing sensing efficiency with limited budgets. The proposed SPO framework integrates a constrained QP optimization layer in the neural network and derives an explicit alternating differentiation method based on the ADMM. The matrix-based formulation in optimization enables seamless integration of the optimization layer into deep learning frameworks, facilitating efficient training and deployment. The unrolling alternating differentiation approach within the optimization layer enables rapid and effective backpropagation, essential for training deep learning models.
The effectiveness of the proposed SPO framework is validated through real-world experiments conducted in both mid-size and large-scale networks within the Kowloon district, Hong Kong. Results demonstrate that the SPO framework with alternating differentiation method surpasses benchmarks set by the SPO with implicit differentiation techniques and PTO in large-scale networks. Sensitivity analyses on target distribution uncertainties and control ratios validate the robustness of the framework for various applications. It is worth noting that the SPO framework (including both SPO-A and SPO-C) achieves a relatively higher prediction error than the two-step PTO framework in most cases, while also yielding better matching accuracy.
Future directions of this research could focus on exploring further explicit differentiation methods specifically tailored for optimization layers, aimed at significantly enhancing computational efficiency, particularly in large-scale networks. Additionally, we aim to extend the application of the SPO framework to encompass other critical traffic management scenarios, such as vehicle routing, which presents an opportunity to tackle more intricate real-world challenges effectively. These advancements would not only expand the applicability of the SPO framework but also contribute to advancing the state-of-the-art in intelligent transportation systems.
Acknowledgments
The work described in this paper is supported by grants from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project No. PolyU/25209221 and PolyU/15206322) and a grant from the Otto Poon Charitable Foundation Smart Cities Research Institute (SCRI) at the Hong Kong Polytechnic University (Project No. P0043552). The contents of this article reflect the views of the authors, who are responsible for the facts and accuracy of the information presented herein.
Appendix A Notations
Table 6 presents the notations in the paper and Table 7 summarizes the dimensions of the vectorized variables for the SPO framework.
Table 6: List of notations.
$I$ The set of origin grids
$J$ The set of destination grids, $|I|=|J|=N$.
$T$ The set of all time intervals.
$V$ The set of vehicle class.
Indices
$v$ The index of vehicle type, $v=\{a,c,f\}$. $a$ represents all vehicles, $c$ represents controllable DVs, $f$ represents non-dedicated free vehicles.
$\tau$ The index of the time interval.
$k$ The index of iterations.
$n$ The index of constraints.
Variables as scalars
$D_{v,i}^{\tau}$ The demand of vehicle type $v$ in grid $i$ at time $\tau$.
$A_{f,i}^{\tau}$ The adjacent information for NDVs in grid $i$ at time $\tau$.
$H_{f,i}^{\tau}$ The hidden information for NDVs in grid $i$ at time $\tau$.
$x_{v,ij}^{\tau}$ The vehicle flow for vehicle type v from origin $i$ to destination $j$ at time $\tau$.
$w_{c,ij}^{\tau}$ The incentive cost from origin $i$ to destination $j$ for DVs at time $\tau$.
$m_{c,ij}^{\tau}$ The travel time from origin $i$ to destination $j$ for DVs at time $\tau$.
Variables/Parameters as tensors
$\mathbf{y}_{k}$ The one-dimension vector of vehicle flow in the $k^{th}$ iteration.
$\mathbf{D}_{v}^{\tau}$ The spatial distribution for vehicle type $v$ at time $\tau$.
$\mathbf{T}_{a}^{\tau}$ The target distribution of all vehicles $a$ at time $\tau$.
$\mathbf{s}_{k}$ The slack variables in $k^{th}$ iteration.
$\mathbf{\mu}_{k}$ The dual variables in $k^{th}$ iteration.
Parameters
$\mathscr{L}$ The loss function. $\mathscr{L}_{1}$ is the prediction loss, $\mathscr{L}_{2}$ is the matching loss, and $\mathscr{L}_{SPO}$ is the SPO loss.
$w$ The weight. $w_{1}$ is the weight of the prediction loss in the SPO function, and $w_{2}$ is the weight of the matching loss in the SPO framework.
$\mathbf{w}_{p}$ The weight in the prediction model.
$\delta$ The time length for each time interval.
$\alpha$ The Look back window in the prediction module.
$\mathbf{C}$ The incentive cost vector.
$R$ The total budget of the incentive cost.
$K$ The total number of alternating layers.
$\xi$ The convergence threshold.
$\gamma$ The control ratio.
$\rho$ The penalty term.
Table 7: Dimension of the vectors in the SPO framework.
Vector Dimension Type Vector Dimension Type
Variables
$\mathbf{y}$ $\mathbb{R}^{N^{2}}$ Vector
Parameters
$\mathbf{A}$ $\mathbb{R}^{N\times N^{2}}$ Sparse Matrix $\mathbf{B}$ $\mathbb{R}^{N\times N^{2}}$ Sparse Matrix
$\mathbf{C}$ $\mathbb{R}^{N^{2}}$ Vector ${\mathbf{D}_{v}^{\tau}}$ $\mathbb{R}^{N}$ Vector
$\mathbf{P}$ $\mathbb{R}^{N^{2}\times N^{2}}$ Sparse Matrix $\mathbf{q}$ $\mathbb{R}^{N^{2}}$ Vector
$\mathbf{G_{1}}$ $\mathbb{R}^{N\times N^{2}}$ Sparse Matrix $\mathbf{s_{1}}$ $\mathbb{R}^{N}$ Vector
$\mathbf{G_{2}}$ $\mathbb{R}^{N^{2}\times N^{2}}$ Sparse Matrix $\mathbf{s_{2}}$ $\mathbb{R}^{N^{2}}$ Vector
$\mathbf{G_{3}}$ $\mathbb{R}^{1\times N^{2}}$ Vector $\mathbf{s_{3}}$ $\mathbb{R}^{1}$ Constant
$\mathbf{G_{4}}$ $\mathbb{R}^{N^{2}\times N^{2}}$ Sparse Matrix $\mathbf{s_{4}}$ $\mathbb{R}^{N^{2}}$ Vector
$\mathbf{h_{1}}$ $\mathbb{R}^{N}$ Vector $\mathbf{\mu_{1}}$ $\mathbb{R}^{N}$ Vector
$\mathbf{h_{2}}$ $\mathbb{R}^{N^{2}}$ Vector $\mathbf{\mu_{2}}$ $\mathbb{R}^{N^{2}}$ Vector
$\mathbf{h_{3}}$ $\mathbb{R}^{1}$ Vector $\mathbf{\mu_{3}}$ $\mathbb{R}^{1}$ Constant
$\mathbf{h_{4}}$ $\mathbb{R}^{N^{2}}$ Vector $\mathbf{\mu_{4}}$ $\mathbb{R}^{N^{2}}$ Vector