### v3

# SPO-VCS: An End-to-End Smart Predict-then-Optimize Framework with Alternating Differentiation Method for Relocation Problems in Large-Scale Vehicle Crowd Sensing

Abstract
Ubiquitous mobile devices have catalyzed the development of vehicle crowd sensing (VCS). In particular, vehicle sensing systems show great potential in the flexible acquisition of extensive spatio-temporal urban data through built-in smart sensors under diverse sensing scenarios. However, vehicle systems like taxis often exhibit biased coverage due to the heterogeneous nature of trip requests and varying routes. To achieve a high sensing coverage, a critical challenge lies in how to optimally relocate vehicles to minimize the divergence between the spatio-temporal distributions of vehicles and target sensing distributions. Conventional approaches typically employ a two-stage predict-then-optimize (PTO) process: first predicting real-time vehicle distributions and subsequently generating an optimal relocation strategy based on the predictions. However, this approach can lead to suboptimal decision-making due to the propagation of errors from upstream prediction. To this end, we develop an end-to-end Smart Predict-then-Optimize (SPO) framework by integrating optimization into prediction within the deep learning architecture, and the entire framework is trained by minimizing the task-specific matching divergence rather than the upstream prediction error. Methodologically, we formulate the vehicle relocation problem by quadratic programming (QP) and incorporate a novel unrolling approach based on the Alternating Direction Method of Multipliers (ADMM) within the SPO framework to compute gradients of the QP layer, facilitating backpropagation and gradient-based optimization for end-to-end learning. The effectiveness of the proposed framework is validated using two real-world taxi datasets ranging from mid-size to large-scale in Hong Kong and Chengdu, China. Utilizing the alternating differentiation method, the general SPO framework presents a novel concept of addressing decision-making problems with uncertainty, demonstrating significant potential for advancing applications in logistics and intelligent transportation systems.
keywords:
Vehicle Relocation , Vehicle Crowd Sensing (VCS) , Smart Predict-then-Optimize (SPO) , Alternating Direction Method of Multipliers (ADMM) , Unrolling Approach
journal: Transportation Research Part E: Logistics and Transportation Review
1 Introduction
In recent years, the development of ubiquitous mobile devices such as mobile phones has significantly advanced the field of mobile crowd sensing (MCS) (Ganti et al. 2011). Compared with the traditional fixed sensor networks, MCS offers more flexible and extensive coverage of sensing information while reducing the installation and collection cost (Li et al. 2018a). Individuals participating in MCS utilize the built-in smart sensors in mobile devices such as cameras and GPS, to sense the surroundings and gather city-wide information (Li et al. 2022b).
The MCS can be categorized into community sensing (CS) and vehicle crowd sensing (VCS) based on different types of sensor hosts (Ji et al. 2023a). CS utilizes sensors installed in mobile devices, while VCS leverages the sensing capabilities of vehicles to collect city-wide data (Xu et al. 2019). Host vehicles of VCS include taxis (Honicky et al. 2008), trams (Saukh et al. 2014), buses (Kang et al. 2016), and unmanned aerial vehicles (UAVs) (Li et al. 2022a). In general, the advantages of VCS in terms of temporal duration, geographical coverage, reliability, and consistency make it particularly suitable for scenarios requiring continuous, full coverage, and high-quality data (Ji et al. 2023a).
Consequently, VCS is instrumental in domains that require high-precision data, such as infrastructure monitoring (Song et al. 2020, Hasenfratz et al. 2015, Liu and Ma 2022), traffic state estimation (Guo et al. 2022, Zhu et al. 2014) and traffic management (Ji et al. 2023b, Dai and Han 2023).
Specifically, the vehicles in the VCS framework are classified into two types: dedicated vehicles (DVs) and non-dedicated vehicles (NDVs). DVs, such as vehicles equipped with air pollution sensors and delivery drones with cameras, are fully guided by the dispatching center to prioritize sensing tasks. In contrast, NDVs, such as e-hailing vehicles and taxis, are under the complete control of their drivers and do not follow the guidance of the dispatching center. Upon receiving a sensing task, DVs will proceed immediately to the specific sensing locations. In contrast, NDVs operate under the control of drivers and are confined to their preferred cruising areas and routes. Moreover, NDVs can be converted into DVs through incentives such as monetary compensation.
Figure 1 provides an overview of the sensing procedure in the VCS framework. The data request platform initially publishes sensing tasks. Upon receiving these tasks, the dispatching center develops relocation strategies for DVs. During the subsequent time intervals, all the sensing vehicles, including DVs and NDVs, move to the specified destinations to perform the assigned sensing tasks. After collecting the necessary data, the sensing vehicles will transmit sensing information back to the data request platform.
Figure 1: An illustration of the sensing procedure in the VCS framework.
In the domain of VCS, the target distribution of sensing tasks differs markedly from the spatio-temporal distribution of sensors. Target distributions typically vary based on specific sensing objectives and adapt over time and space. For instance, monitoring air quality may necessitate uniformly distributed data across a city (Bales et al. 2012, Chen et al. 2018), whereas tasks focusing on industrial pollution may require the spatial distribution in Gaussian (Khedo et al. 2010, Paulos et al. 2007). In contrast, the spatio-temporal distributions of sensors are heavily influenced by the movements of host sensing vehicles (Anjomshoaa et al. 2018).
Therefore, a critical challenge in VCS lies in the effective relocation of vehicles to align with the target sensing distribution. To be specific, the dispatching center needs to develop relocation strategies for DVs to minimize the divergence between the target sensing distributions and vehicle distributions. The state-of-the-art approach to addressing the vehicle relocation problem in VCS is a two-stage predict-then-optimize (PTO) method, where a prediction model is first established to forecast the movements of NDVs (Chang et al. 2023, Simchi-Levi 2014, Mišić and Perakis 2020). The prediction is then used as the input for the downstream optimization problem related to DVs. However, a key limitation of the PTO approach is that training the prediction model is mainly based on empirical error minimization, and metrics such as Mean Squared Error (MSE) and Mean Absolute Error (MAE) are used. The trained prediction model may not necessarily yield optimal strategies in the subsequent optimization stage, because the focus of the optimization stage is task-specific, which may not align well with the metrics of MSE or MAE (Bengio 1997, Ford et al. 2015, Elmachtoub and Grigas 2022, Yan and Wang 2022). This discrepancy highlights the need for more integrated approaches that bridge prediction and optimization phases coherently in the context of VCS.
Recently, the end-to-end Smart Predict-then-Optimize (SPO) framework has been proposed and attracted wide attention (Elmachtoub and Grigas 2022). This framework integrates prediction and optimization within a unified deep-learning architecture. Specifically, the optimization module operates as a differentiation layer, mapping inputs directly to optimal solutions without requiring explicit closed-form solutions. This approach enables the entire neural network to be trained by directly minimizing task-specific errors. However, the application of the SPO framework to address challenges such as vehicle relocation has not been extensively explored.
The main challenge in applying the SPO framework is integrating the optimization layer seamlessly within the deep learning architecture. A widely used approach is the implicit differentiation, where the Karush-Kuhn-Tucker (KKT) conditions of the optimization problems are differentiated to obtain first-order derivatives of the optimal solution with respect to the parameters. However, this method can be computationally demanding in directly computing the Jacobian matrix of the KKT conditions for large-scale problems (Diamond and Boyd 2016, Agrawal et al. 2019, Tang and Khalil 2022).
In this paper, we advocate an alternative approach - explicit unrolling, which avoid the direct computation of the Jacobian matrix in the KKT conditions and compute the corresponding gradients iteratively, enhancing computational efficiency and suitability for large-scale problems (Sun et al. 2023, Li et al. 2020). Nevertheless, the majority of unrolling approaches are predominantly applicable to unconstrained problems. In the task of VCS, relocation challenges are usually formulated as constrained optimization problems. To this end, we incorporate the Alternating Direction Method of Multipliers (ADMM) to solve constrained optimization problems by decomposing the original problem into a series of unconstrained sub-problems (Boyd et al. 2011). Combining the ADMM with differentiation methods and unrolling approaches can potentially address the challenge of integrating large-scale constrained convex optimization problems into deep learning architectures.
To conclude, there are two major issues with the predict-and-optimize problem in VCS. First, the conventional PTO framework for vehicle relocation problems may not generate an optimal relocation strategy, resulting in a substantial divergence between the desired target distribution and the actual vehicle distribution. Second, embedding an optimization layer within neural networks is usually computationally intensive and impractical for large-scale networks, and existing unrolling approaches are often inadequate for handling constrained optimization problems effectively.
To address these challenges, we develop an innovative SPO framework coupled with an alternating differentiation method for vehicle relocation problems in VCS, aiming to achieve optimal sensing coverage. Specifically, we formulate the vehicle relocation problem based on constrained quadratic programming (QP) and embed the QP layer into neural networks. To our knowledge, this is the first time that the SPO framework coupled with an alternating differentiation method has been applied in vehicle relocation in mobile sensing. Our contributions are outlined as follows:
  • 1.
    A novel end-to-end Smart Predict-then-Optimize framework for vehicle relocation problems. We first time formulate the vehicle relocation problem using an SPO framework in deep learning. Specifically, the vehicle relocation problem is formulated as a constrained QP, where the parameters in its objectives are predicted using deep learning.
  • 2.
    An alternating differentiation approach based on the ADMM for constrained optimization problems in large-scale networks. A novel unrolling approach is proposed to obtain the gradients of the constrained QP iteratively for large-scale networks. Specifically, the forward pass is solved using the ADMM, and the backpropagation is achieved through alternating differentiation.
  • 3.
    Two large-scale real-world experiments. Numerical experiments on two real-world datasets across five network sizes from mid-size to large-scale are conducted. Superior matching performance and computational efficiency of the proposed SPO framework are demonstrated.
The remainder of the paper is organized as follows: Section 2 reviews the related work on the vehicle relocation problem in mobile sensing. Section 3 introduces the SPO framework in vehicle relocation, including the prediction module, the optimization module, the integrated model, and an unrolling approach based on the alternating differentiation method for the SPO framework. Section 4 presents the overall solution algorithm of the end-to-end SPO framework. Section 5 evaluates the proposed framework in two real-world scenarios and compares the method with the baseline approaches. Section 6 concludes the results and proposes future directions.
2 Literature Review
This section provides a comprehensive review of the literature related to the vehicle relocation problem in mobile sensing, the end-to-end SPO framework and its differentiation approach, and applications of the SPO framework in the field of transportation.
2.1 Vehicle relocation in mobile sensing
Pervasive crowd sensing plays an important role in intelligent transport systems, logistics and smart city applications (Ji et al. 2023a, Dai and Han 2023, Yang et al. 2025a). Given their relatively lower installation costs and extended operational duration, taxis are the most commonly utilized hosts in the VCS. However, due to heterogeneous trip patterns, the spatio-temporal distributions of these vehicles are often biased. To enhance sensing quality, numerous scholars have investigated intervention measures to effectively relocate vehicles through the development of incentivizing schemes (Fan et al. 2019, Xu et al. 2019, Chen et al. 2020) or proactive scheduling (Masutani 2015).
Existing research on solving the conventional vehicle relocation problem can be categorized into three types: the steady-state method, the model-free method, and the real-time optimization method (Lei et al. 2020, Qian et al. 2022). The steady-state method modeled the vehicle relocation system through a queuing theoretical framework and solved the optimal relocation strategy when the system reaches a steady state (Pavone et al. 2012, Sayarshad and Chow 2017, Chuah et al. 2018). A significant limitation of this approach is the inadequate capacity to manage the complexities of large-scale problems. In contrast, the model-free method, based on deep reinforcement learning (DRL), has emerged as a promising approach for its merits in addressing large-scale vehicle relocation challenges (Qian et al. 2022, Jiao et al. 2021, Shou and Di 2020, Qin et al. 2022).
Another classical approach in vehicle sensing is the real-time optimization method, which formulates the vehicle relocation problem from an optimization perspective and identifies optimal strategies through various solutions (Kek et al. 2009, Huang et al. 2018). Miao et al. (2016) and Zhang et al. (2016) formulated the vehicle scheduling and routing problem as a mixed integer program (MIP) and utilized the model predictive control (MPC) method to solve it. Recent advancements have also incorporated deep learning techniques to forecast future travel patterns, employing a two-stage PTO framework for the vehicle relocation problem (Weikl and Bogenberger 2013, Chang et al. 2022; 2023). This method first estimates the spatio-temporal vehicle patterns using deep learning, followed by solving the downstream optimization problem with the prediction results as inputs. However, a notable limitation of the PTO approach is that training the prediction model solely based on prediction error can result in suboptimal decision-making, as opposed to directly minimizing the decision error (Agrawal et al. 2019, Elmachtoub and Grigas 2022). Even with optimal predictions yielding minimal prediction errors, the resulting optimization may not be optimal. Our study extends the real-time PTO framework by integrating optimization directly within the prediction process through a deep learning architecture.
Table 1 summarizes related works of the vehicle relocation problem in mobile sensing and their fundamental settings. In a nutshell, our work differs from the previous research in three key aspects. First, unlike most existing literature that fully controls vehicles, we classify vehicles into dedicated and non-dedicated types, allowing the non-dedicated ones to operate freely. Second, to our knowledge, we are the first to apply the SPO framework to vehicle relocation in mobile sensing and validate it on large-scale networks. Third, we evaluate the framework across multiple target distributions, an aspect rarely explored in previous studies.
Table 1: The related work of vehicle relocation in mobile sensing.
Acknowledgments
The work described in this paper is supported by grants from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project No. PolyU/25209221 and PolyU/15206322) and a grant from the Otto Poon Charitable Foundation Smart Cities Research Institute (SCRI) at the Hong Kong Polytechnic University (Project No. P0043552). The contents of this article reflect the views of the authors, who are responsible for the facts and accuracy of the information presented herein.
Appendix A Notations
Table A1 presents the notations in the paper, and Table A2 summarizes the dimensions of the vectorized variables for the SPO framework.
Table A1: List of notations.
II The set of origin grids
$J$ The set of destination grids, $|I|=|J|=N$.
$T$ The set of all time intervals.
$V$ The set of vehicle class.
Indices
$v$ The index of vehicle type, $v=\{a,c,f\}$. $a$ represents all vehicles, $c$ represents controllable DVs, $f$ represents non-dedicated free vehicles.
$\tau$ The index of the time interval.
$k$ The index of iterations.
$n$ The index of constraints.
Variables as scalars
$D_{v,i}^{\tau}$ The demand of vehicle type $v$ in grid $i$ at time $\tau$.
$A_{f,i}^{\tau}$ The adjacent information for NDVs in grid $i$ at time $\tau$.
$H_{f,i}^{\tau}$ The hidden information for NDVs in grid $i$ at time $\tau$.
$x_{v,ij}^{\tau}$ The vehicle flow for vehicle type v from origin $i$ to destination $j$ at time $\tau$.
$w_{c,ij}^{\tau}$ The incentive cost from origin $i$ to destination $j$ for DVs at time $\tau$.
$m_{c,ij}^{\tau}$ The travel time from origin $i$ to destination $j$ for DVs at time $\tau$.
Variables/Parameters as tensors
$\mathbf{y}_{k}$ The one-dimension vector of vehicle flow in the $k^{th}$ iteration.
$\mathbf{D}_{v}^{\tau}$ The spatial distribution for vehicle type $v$ at time $\tau$.
$\mathbf{T}_{a}^{\tau}$ The target distribution of all vehicles $a$ at time $\tau$.
$\mathbf{s}_{k}$ The slack variables in $k^{th}$ iteration.
$\mathbf{\mu}_{k}$ The dual variables in $k^{th}$ iteration.
Parameters
$\mathscr{L}$ The loss function. $\mathscr{L}_{1}$ is the prediction loss, $\mathscr{L}_{2}$ is the matching loss, and $\mathscr{L}_{SPO}$ is the SPO loss.
$w$ The weight. $w_{1}$ is the weight of the prediction loss in the SPO function, and $w_{2}$ is the weight of the matching loss in the SPO framework.
$\mathbf{w}_{p}$ The weight in the prediction model.
$\delta$ The time length for each time interval.
$\alpha$ The Look back window in the prediction module.
$\mathbf{C}$ The incentive cost vector.
$R$ The total budget of the incentive cost.
$K$ The total number of alternating layers.
$\xi$ The convergence threshold.
$\gamma$ The control ratio.
$\rho$ The penalty term.
Table A2: Dimension of the vectors in the SPO framework.
Vector Dimension Type Vector Dimension Type
Variables
$\mathbf{y}$ $\mathbb{R}^{N^{2}}$ Vector
Parameters
$\mathbf{A}$ $\mathbb{R}^{N\times N^{2}}$ Sparse Matrix $\mathbf{B}$ $\mathbb{R}^{N\times N^{2}}$ Sparse Matrix
$\mathbf{C}$ $\mathbb{R}^{N^{2}}$ Vector $\mathbf{D}_{v}^{\tau}$} $\mathbb{R}^{N}$ Vector
$\mathbf{P}$ $\mathbb{R}^{N^{2}\times N^{2}}$ Sparse Matrix $\mathbf{q}$ $\mathbb{R}^{N^{2}}$ Vector
$\mathbf{G_{1}}$ $\mathbb{R}^{N\times N^{2}}$ Sparse Matrix $\mathbf{s_{1}}$ $\mathbb{R}^{N}$ Vector
$\mathbf{G_{2}}$ $\mathbb{R}^{N^{2}\times N^{2}}$ Sparse Matrix $\mathbf{s_{2}}$ $\mathbb{R}^{N^{2}}$ Vector
$\mathbf{G_{3}}$ $\mathbb{R}^{1\times N^{2}}$ Vector $\mathbf{s_{3}}$ $\mathbb{R}^{1}$ Constant
$\mathbf{G_{4}}$ $\mathbb{R}^{N^{2}\times N^{2}}$ Sparse Matrix $\mathbf{s_{4}}$ $\mathbb{R}^{N^{2}}$ Vector
$\mathbf{h_{1}}$ $\mathbb{R}^{N}$ Vector $\mathbf{\mu_{1}}$ $\mathbb{R}^{N}$ Vector
$\mathbf{h_{2}}$ $\mathbb{R}^{N^{2}}$ Vector $\mathbf{\mu_{2}}$ $\mathbb{R}^{N^{2}}$ Vector
$\mathbf{h_{3}}$ $\mathbb{R}^{1}$ Vector $\mathbf{\mu_{3}}$ $\mathbb{R}^{1}$ Constant
$\mathbf{h_{4}}$ $\mathbb{R}^{N^{2}}$ Vector $\mathbf{\mu_{4}}$ $\mathbb{R}^{N^{2}}$ Vector
Appendix B Hyper-parameters in different prediction modules.
Table B1 lists the hyper-parameters used in various prediction models.
Table B1: Hyper-parameters of different prediction modules.
   Model    Hyper-parameter    Value
   TGCN    Number of hidden layers    2
   Number of channels in each layer    64, 32
   Kernel size    2
   Dropout    0.2
   Activation function    ReLU
   DCRNN    Number of hidden layers    2
   Number of channels in each layer    64, 32
   Dropout    0.2
   Activation function    ReLU
   ASTGCN    Number of hidden layers    2
   Number of channels in each layer    32, 16
   Number of heads    2
   Dropout    0.2
   Activation function    ReLU